If an alkaline earth metal combines with a chalcogen to form 7.88 * 10^23 molecules how many atoms are involved?
1
Expert's answer
2020-02-28T12:13:21-0500

Li + O 2  = 2LiO

4 * 6.02 * 1023 + 1 * 6.02 * 1023 = 24.08 * 1023 + 6.02 * 10 23 = 30.1 * 1023